Green light of wavelength 550 nm is shone on potassium, which

has a work function of 2.0 eV. Calculate the maximum kinetic energy of the photoelectrons emitted.


1
Expert's answer
2021-01-20T08:42:56-0500

We can find the maximum kinetic energy of the photoelectrons emitted from the formula:


"KE_{max}=hf-W,""KE_{max}=\dfrac{hc}{\lambda}-W,""KE_{max}=\dfrac{6.63\cdot10^{-34}\ J\cdot s\cdot3\cdot10^8\ \dfrac{m}{s}}{550\cdot10^{-9}\ m}-2.0\ eV,""KE_{max}=3.62\cdot10^{-19}\ J\cdot\dfrac{1\ eV}{1.6\cdot10^{-19}\ J}-2.0\ eV=0.26\ eV."

Answer:

"KE_{max}=0.26\ eV."